# The nature of mathematical models

### v1

@@ 3.3 Linear and nonlinear models
Denoting $\dot{\boldsymbol{X}}:=\boldsymbol{X}^{\prime}$ we have that $(\boldsymbol{X}^{i})^{\prime}=(\dot{\boldsymbol{X}})^{i}=\dot{\boldsymbol{X}}^{i}$     hence the tangent space is the span $T_{\hat{\boldsymbol{X}}}{\mathcal{M}}=\mathcal{S}\{\dot{\boldsymbol{X}}^{i}\}_{i=1,...,q}$    WRT the canonical model chart map $(\boldsymbol{X}^{-1})$. This means that for any $\eta\in\mathbb{R}^{q}$, $(\eta\cdot\dot{\boldsymbol{X}}+\hat{\boldsymbol{X}})\in T_{\hat{\boldsymbol{X}}}\mathcal{M}\subset\mathcal{H}$ and $(\eta\cdot\dot{\boldsymbol{X}}+\hat{\boldsymbol{X}})$ is a random variable in the (affine) tangent space to the model manifold $\mathcal{M}$ at $\hat{\boldsymbol{X}}$.
An affine subspace $T_{\hat{\boldsymbol{X}}}\mathcal{M}\subset\mathcal{H}$ of random variables has thus been constructed using the model canonical chart map $\boldsymbol{X}^{-1}$. The Hilbert projection theorem applies and the random variable $\boldsymbol{X}^{o}$ has a unique projection onto $\hat{\boldsymbol{X}}\in T_{\hat{\boldsymbol{X}}}\mathcal{M}$ , of global minimal distance from $T_{\hat{\boldsymbol{X}}}\mathcal{M}$, such that $\forall\,\boldsymbol{X}\in T_{\hat{\boldsymbol{X}}}\mathcal{M}$ we have $(\boldsymbol{X}^{o}-\hat{\boldsymbol{X}})\,\perp(\boldsymbol{X}-\hat{\boldsymbol{X}})$.
@@ 3.5 Relationship with Information Geometry
The well-developed field of Information Geometry deals with spaces of distributions, wherein certain (parametrized) families are embedded manifolds. Interesting properties can be derived for some such families (e.g. the exponential family generates a "flat" manifold).
Very interesting facts emerge from the study of such spaces of distributions. For example, denote by $\mathcal{P}$ the space of probability distributions, and denote by $\mathcal{F}\subset\mathcal{P}$ the manifold representing a family of probability distributions indexed by some parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{q}$, $F(\boldsymbol{\theta})\in\mathcal{F}$, which family we wish to use as a possible “model” for our data-generating process. We assume that the data are actually generated by an unknown distribution $P\in\mathcal{P}$, with, in general $P\notin\mathcal{F}$. If we observe data $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ we can build an empirical distribution $\tilde{P}$ out of them. We could then hypothesize that the “best” value of $\boldsymbol{\theta}$ is the value $\hat{\boldsymbol{\theta}}$ that minimizes some kind of distance or divergence from $\tilde{P}$ to $F(\boldsymbol{\theta})$: in the case the Kullback-Leibler divergence $D_{KL}$ of $F(\boldsymbol{\theta})$ from $\tilde{P}$ is considered, we obtain the Maximum Likelihood Estimator $\hat{\boldsymbol{\theta}}_{ML}$.
@@
3.6 Mapping $\mathcal{H}$ to $\mathbb{R}^{n}$
We need now to formalize the intuitively obvious relationship between the theoretical construction of the abstract model in the Hilbert space of finite-variance random variables $\mathcal{H}$ and the actual computational, numerical analysis of the results carried out with statistical techniques in case space $\mathbb{R}^{n}$. A random variable is defined as a measurable function from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{n}$, say
$Y\in\mathcal{H}\quad\Rightarrow\quad Y:\Omega\rightarrow\mathbb{R}^{n}\,\,,\,\,\,Y:\omega\mapsto r\,.$
However, we might equally well define an event $\omega\in\Omega$ as a map from $\mathcal{H}$ to $\mathbb{R}^{n}$:
$\omega\in\Omega\quad\Rightarrow\quad\omega:\mathcal{H}\rightarrow\mathbb{R}^{n}\,\,,\,\,\,\omega:Y\mapsto r\,.$
In other words, we may define
$\omega(Y):=Y(\omega)=r\in\mathbb{R}^{n}$ (4)
Experimental sampling is the concrete act of picking an event $\omega$ from the sample space $\Omega$. This act determines the specific realization $\boldsymbol{x}^{o}$ of the random variable $\boldsymbol{X}^{o}$. As a consequence, this act also determines the realization $\hat{\boldsymbol{x}}$ of $\hat{\boldsymbol{X}}$, and so on. All random variables considered in the theoretical construction of the model are thus naturally mapped by $\omega$ to their respective numerical realizations. We see then that the event $\omega$ can be interpreted as the necessary map from $\mathcal{H}$ to $\mathbb{R}^{n}$, see Fig.4.
Figure 4: Sampling as a map from $\mathcal{H}$ to $\mathbb{R}^{n}$.
4 Discussion and Conclusions
@@
Recent literature addresses similar problems (in particular the characterization of the projection) in the case of conditional expectation [4] and in the case of information geometry [6, 7, 10, 8, 9]. Specifically, in information geometry a set of probability distributions indexed by a finite number of parameters is shown to be a (smooth) manifold and is equipped with a Riemannian structure and with conjugate connections, which make the manifold ‘flat’ and generalize the concept of Euclidean Space [8, 6]. This ‘flat’ structure can be obtained in various ways, some of which are discussed in the papers of Amari et al. and Nielsen [9, 8, 6, 11]. One way of doing this is through Bregman divergence, which induces a structure on a manifold $\mathcal{M}$ that generalizes the Pythagorean Theorem and consequently guarantees the existence and uniqueness of the projection of a point $P$ onto the manifold $\mathcal{M}$ [9, 8]. A similar approach to Amari (also formulating a projection theorem) is proposed by Brigo [10] using a direct $\mathcal{L}^{2}$ metric. We note that, as stated above, the theory of conditional expectation as a projection onto an appropriate subspace of the Hilbert space $\mathcal{H}$ of random variables is not related to the current discussion because the underlying manifold is different. As regards information geometry, there is a clear connection since any random variable induces a probability distribution, but the identification of the underlying manifold as determined by the mathematical model of the (biological) experiment, as well as the identification of the map between the space $\mathcal{H}$, where the modeling is abstractly done, and $\mathbb{R}^{n}$, where the computations are carried out, to the best of our knowledge have not yet been formalized.
@@
In conclusion, a mathematical model is an operator yielding a random variable $\boldsymbol{X}$ approximating the objective random variable $\boldsymbol{X}^{o}$ in $\mathcal{H}$. The event $\omega$ maps $\mathcal{H}$ onto case space $\mathbb{R}^{n}$. Estimation for linear models in case space consists in projecting the observed value $\boldsymbol{x}^{o}$ onto the prediction surface $\mathcal{S}$, and corresponds via $\omega$ to projecting $\boldsymbol{X}^{o}$ onto the model manifold $\mathcal{M}$. For nonlinear models a local optimum $\hat{\boldsymbol{x}}\in\mathcal{S}$ can be found numerically, and is the image under $\omega$ of a locally optimal random variable $\hat{\boldsymbol{X}}\in\mathcal{M}$. Asymptotic confidence regions can be computed on $\mathcal{S}$ for linear models. For nonlinear models approximate asymptotic confidence regions can be computed on the tangent space $T_{\hat{\boldsymbol{x}}}(\mathcal{S})$ (constructed via computation of the Jacobian at the local optimum), which corresponds to the theoretical tangent space of differential operators $T_{\hat{X}}(\mathcal{M})\subset\mathcal{H}$.

### v2

@@ 3.3 Linear and nonlinear models
Denoting $\dot{\boldsymbol{X}}:=\boldsymbol{X}^{\prime}$ and writing $(\boldsymbol{X}^{j})^{\prime}=(\dot{\boldsymbol{X}})^{j}=\dot{\boldsymbol{X}}^{j}$  ,   the tangent space is the span $T_{\hat{\boldsymbol{X}}}{{\mathcal{M}_{\boldsymbol{X}}}}={\mathcal{S}}\{\dot{\boldsymbol{X}}^{j}\}_{j=1,...,q}$    WRT the canonical model chart map $(\boldsymbol{X}^{-1})$. This means that for any ${\boldsymbol{\eta}}\in\mathbb{R}^{q}$, $\dot{\boldsymbol{X}}\cdot{\boldsymbol{\eta}}\in T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}$ and that $\mathcal{T}_{\hat{{\boldsymbol{X}}}}=\hat{\boldsymbol{X}}+T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}=\hat{\boldsymbol{X}}+\{\dot{\boldsymbol{X}}\cdot{\boldsymbol{\eta}}\,,\,\,{\boldsymbol{\eta}}\in\mathbb{R}^{q}\}\subset\mathcal{H}$ is an affine space of random variables tangent to the model manifold $\mathcal{M}_{\boldsymbol{X}}$} at $\hat{\boldsymbol{X}}$.
[…]
3.4 Mapping $\mathcal{H}$ to $\mathbb{R}^{n}$
It is now possible to formalize the intuitively obvious relationship between the theoretical construction of the abstract model in the Hilbert space of finite-variance random variables $\mathcal{H}$ and the actual computational, numerical analysis of the results carried out with statistical techniques in case space $\mathbb{R}^{n}$.
Recall that a random variable is defined as a measurable function from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{n}$, mapping outcomes to (vectors of) reals, say
${\boldsymbol{Y}}\in\mathcal{H}\quad\Rightarrow\quad{\boldsymbol{Y}}:\Omega\rightarrow\mathbb{R}^{n}\,\,,\,\,\,{\boldsymbol{Y}}:\omega\mapsto r\,.$
In fact, the act of sampling some variable on some experimental unit (a person from a population of interest, the roll of a die, …) can be seen as a map from the product space $\Omega\times{\mathcal{H}}$ to $\mathbb{R}^{n}$, of which ${\boldsymbol{Y}}:\Omega\rightarrow\mathbb{R}^{n}$ is the marginal map obtained by fixing $\boldsymbol{Y}$$\in{\mathcal{H}}$. In the same way we may consider the other marginal map, obtained by fixing $\omega\in\Omega$: we can thus consider the outcome as a map from $\mathcal{H}$ to $\mathbb{R}^{n}$:
$\omega\in\Omega\quad\Rightarrow\quad\omega:\mathcal{H}\rightarrow\mathbb{R}^{n}\,\,,\,\,\,\omega:{\boldsymbol{Y}}\mapsto r\,.$
In other words, we may define
$\omega({\boldsymbol{Y}}):={\boldsymbol{Y}}(\omega)=\boldsymbol{r}\in\mathbb{R}^{n}$ (4)
Experimental sampling is the concrete act of picking an outcome $\omega$ from the sample space $\Omega$. This act fixes $\omega$ and thus determines the specific realizations $\boldsymbol{x}^{o},\,\boldsymbol{y}^{o},\dots$ of any desired random variables $\boldsymbol{X}^{o},\boldsymbol{Y}^{o},\dots$. As a consequence and within the context of our modeling discussion, the sampling act also determines the realization $\hat{\boldsymbol{x}}$ of $\hat{\boldsymbol{X}}$, and so on. All random variables considered in the theoretical construction of the model are thus naturally mapped by $\omega$ to their respective numerical realizations. We see then that the event $\omega$ can be interpreted as the necessary map from $\mathcal{H}$ to $\mathbb{R}^{n}$, see Fig.4.
Figure 4: Sampling as a map from $\mathcal{H}$ to $\mathbb{R}^{n}$.
3.5 Summary
We recap at this point the different elements of the twin structures (in $\mathcal{H}$ and in $\mathbb{R}^{n}$) that we have described. While the use of many of these elements corresponds to standard statistical practice, the following list has the purpose of systematizing the notation to dispel possible sources of confusion, as well as of clarifying the correspondence of objects in the abstract modeling and the concrete computational settings.
  • We are interested in predicting the values of some objective random variable $\boldsymbol{X}^{o}\in\mathcal{H}$;
  • to do so we formulate a conceptual model as the operator $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U})$, yielding a random variable $\boldsymbol{X}\in\mathcal{H}$ for any given value of $\boldsymbol{\theta}\in\mathbb{R}^{q}$. We hope to find some optimal value of the parameter $\boldsymbol{\theta}$ such that the random variable $\boldsymbol{X}$ is close to the objective random variable $\boldsymbol{X}^{o}$, this would give us a “good” model.
  • $\boldsymbol{x}^{o}$ are measurements of some quantity of interest, $\boldsymbol{x}^{o}\in\mathbb{R}^{n}$. We approximate $\boldsymbol{x}^{o}$ by means of some other quantity $\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{u})=\boldsymbol{x}(\boldsymbol{\theta})$ computed as a function of some parameter $\boldsymbol{\theta}\in\Theta\subset\mathbb{R}^{q}$ (and possibly of some predictor values $\boldsymbol{u}$): the fundamental idea is that we obtain a forecast $\boldsymbol{x}$ for each arbitrary value of $\boldsymbol{\theta}$.
    Notice that we distinguish between a model, i.e. a model operator $\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U}):\mathbb{R}^{q}\times\mathcal{H}\rightarrow\mathcal{H}$, and a model function $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{u}):\mathbb{R}^{q}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, even though their functional form is the same.
    For convenience, disregarding any predictor variable(s) we also define the model map $\boldsymbol{X}(\boldsymbol{\theta}):\mathbb{R}^{q}\rightarrow\mathcal{H}$, which is the inverse of the model-induced (global) canonical chart $\boldsymbol{X}^{-1}:\mathcal{H}\rightarrow\mathbb{R}^{q}$.
  • So $\boldsymbol{\theta}$ is some arbitrary value of the parameter; we believe $\boldsymbol{\theta}^{*}$ to be the "true" but unknown value of the parameter, to which therefore corresponds a true but unknown $\boldsymbol{x}^{*}=\boldsymbol{x}(\boldsymbol{\theta}^{*})$.
  • When making an experiment, we observe the value $\boldsymbol{x}^{o}$, which in general cannot be produced exactly by the function $\boldsymbol{x}(\boldsymbol{\theta})$ for any value of $\boldsymbol{\theta}$; since we have assumed that the relationship is true (with $\boldsymbol{x}^{*}=\boldsymbol{x}(\boldsymbol{\theta}^{*})$), we deduce therefore that some (observation) error has been made in measuring $\boldsymbol{x}^{o}$.
  • We denote this error by $\boldsymbol{\varepsilon}^{o}=\boldsymbol{x}^{o}-\boldsymbol{x}^{*}$ (even though $\boldsymbol{\varepsilon}^{o}$ is not actually observed): we thus understand the observation error that one might make as some random variable $\boldsymbol{\varepsilon}$, of which $\boldsymbol{\varepsilon}^{o}$ is the specific realization obtained in our experiment.
  • We indicate with $\boldsymbol{X}^{o}$ the random variable $\boldsymbol{X}^{o}=\boldsymbol{X}(\boldsymbol{\theta}^{*})+\boldsymbol{\varepsilon}$; $\boldsymbol{x}^{o}$ is the observed realization of $\boldsymbol{X}^{o}$, $\boldsymbol{x}^{o}=\boldsymbol{x}(\boldsymbol{\theta}^{*})+\boldsymbol{\varepsilon}^{o}$, where both $\boldsymbol{\theta}^{*}$ and $\boldsymbol{\varepsilon}^{o}$ are however unknown.
  • The predictors $\boldsymbol{u}$, which we observe, may be regarded as the sample realization of some (design) random variable $\boldsymbol{U}$   (possibly having zero variance, e.g. where fixed by design), hence $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta})=\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U})$, for any given $\boldsymbol{\theta}$, is also a random variable.
  • In order to assess the likely value of the unknown parameter $\boldsymbol{\theta}^{*}$ we introduce an estimator $\hat{\boldsymbol{\theta}}$, which is a function of the random variable $\boldsymbol{X}^{o}$ and hence a random variable itself, $\hat{\boldsymbol{\theta}}=\hat{\boldsymbol{\theta}}(\boldsymbol{X}^{o})$. The statistic $\hat{\boldsymbol{\theta}}^{o}=\hat{\boldsymbol{\theta}}(\boldsymbol{x}^{o})$ is the observed realization of $\hat{\boldsymbol{\theta}}$.
  • $\hat{\boldsymbol{X}}=\boldsymbol{X}(\hat{\boldsymbol{\theta}})=\boldsymbol{X}(\hat{\boldsymbol{\theta}}(\boldsymbol{X}^{o}))$ is the estimator of $\boldsymbol{X}^{o}$ assuming the model $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta})=\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U})$ to be true. Being a function of a random variable, it is of course a random variable itself; its observed realization is $\hat{\boldsymbol{x}}^{o}=\boldsymbol{x}(\hat{\boldsymbol{\theta}}^{o})$.
[…]
  • ${\mathcal{M}_{\boldsymbol{X}}}=\{\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta},\boldsymbol{U}),\boldsymbol{\theta}\in\Theta\}$ is the model manifold, the set of all possible random variables $\boldsymbol{X}$ generated by the model operator. Having observed or set $\boldsymbol{u}$, the prediction surface ${\mathcal{M}_{\boldsymbol{x}}}=\{\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{u}),\boldsymbol{\theta}\in\Theta\}$ is the set of all possible values $\boldsymbol{x}\in\mathbb{R}^{n}$ which can be generated by the model function $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{u})$.
  • $T_{\hat{\boldsymbol{X}}}{{\mathcal{M}_{\boldsymbol{X}}}}$ is the direction space (the tangent space in differential-geometric parlance), the vector space of differential operators defined at $\hat{\boldsymbol{X}}$ with respect to the canonical chart map $\boldsymbol{X}^{-1}$, which is the inverse of the model map $\boldsymbol{X}({\boldsymbol{\theta}})$.
  • $\mathcal{T}_{\hat{{\boldsymbol{X}}}}=\hat{\boldsymbol{X}}+T_{\hat{\boldsymbol{X}}}{\mathcal{M}_{\boldsymbol{X}}}=\hat{\boldsymbol{X}}+\{\dot{\boldsymbol{X}}\cdot{\boldsymbol{\eta}}\,,\,\,{\boldsymbol{\eta}}\in\mathbb{R}^{q}\}\subset\mathcal{H}$ is the affine space of random variables tangent to the model manifold $\mathcal{M}_{\boldsymbol{X}}$} at $\hat{\boldsymbol{X}}$.
  • Notice that $\boldsymbol{x}^{*}=\boldsymbol{x}(\boldsymbol{\theta}^{*},\boldsymbol{u})$ for fixed $\boldsymbol{u}$ is not the realization of a random variable, but it would be if $\boldsymbol{u}$ were a realization of the random variable $\boldsymbol{U}$. In other words, if the design $\boldsymbol{U}$ is fixed, then $\boldsymbol{u}\equiv\boldsymbol{U}$ deterministic and $\boldsymbol{x}^{*}$ is also fixed (but unknown). More generally we could consider $\boldsymbol{x}^{*}$ to be the realization of the random variable $\boldsymbol{X}^{*}=\boldsymbol{x}(\boldsymbol{\theta}^{*},\boldsymbol{U})$, possibly having zero variance.
  • Sometimes we will need to generate samples of artificial realizations of $\boldsymbol{X}$, which we would denote by $\tilde{\boldsymbol{x}}^{r},r=1,2,\dots$.
4 Discussion and Conclusions
Recent literature addresses similar problems (in particular the characterization of the projection) in the case of conditional expectation [Bobrowski2013] and in the case of information geometry [BarndorffNielsen1986, Fletcher2020, Brigo2011, Nielsen2020, Amari2009].
@@ 4.2 Relationship with Information Geometry
The well-developed field of Information Geometry deals with spaces of distributions, wherein certain (parametrized) families are embedded manifolds. Interesting properties can be derived for some such families (e.g. the exponential family generates a “flat” manifold in the appropriate Riemannian metric tensor because, under the Fisher information metric and exponential connection, the parameter space forms an affine, curvature-free geometry).
Interesting facts emerge from the study of such spaces of distributions. For example, denote by $\mathcal{P}$ the space of probability distributions, and denote by $\mathcal{G}\subset\mathcal{P}$ the manifold representing a family of probability distributions indexed by some parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{q}$, $G(\boldsymbol{\theta})\in\mathcal{G}$, which family we wish to use as a possible “model” for our data-generating process. We assume that the data are actually generated by an unknown distribution $P\in\mathcal{P}$, with, in general, $P\notin\mathcal{G}$. If we observe data $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ we can build an empirical distribution $\tilde{P}$ out of them. We could then hypothesize that the “best” value of $\boldsymbol{\theta}$ is the value $\hat{\boldsymbol{\theta}}$ that minimizes some kind of distance or divergence from $\tilde{P}$ to $G(\boldsymbol{\theta})$: in the case the Kullback-Leibler divergence $D_{KL}$ of $G(\boldsymbol{\theta})$ from $\tilde{P}$ is considered, we obtain the Maximum Likelihood Estimator $\hat{\boldsymbol{\theta}}_{ML}$, and if furthermore the family $\mathcal{G}$} is flat (e.g. $\mathcal{G}$} is the exponential family) then estimating by MLE corresponds to projecting $\tilde{P}$ onto $\mathcal{G}$}.
@@
Specifically, in information geometry a set of probability distributions indexed by a finite number of parameters is shown to be a (smooth) manifold and may be equipped with a Riemannian structure and with conjugate connections, which make the manifold ‘flat’, thus generalizing the concept of Euclidean Space [Nielsen2020, BarndorffNielsen1986]. This flat structure can be obtained in different ways, some of which are discussed in the papers of Amari et al. and Nielsen [Amari2009, Nielsen2020, BarndorffNielsen1986, Amari2010]. One way of doing this is through Bregman divergence, which induces a structure on a manifold $\mathcal{G}$} that generalizes the Pythagorean Theorem and consequently guarantees the existence and uniqueness of the projection of a point $\tilde{P}$ onto the manifold $\mathcal{G}$} [Amari2009, Nielsen2020]. A similar approach to Amari (also formulating a projection theorem) is proposed by Brigo [Brigo2011] using a direct $\mathcal{L}^{2}$ metric.
It can therefore be seen that, as stated above, the theory of conditional expectation as a projection onto an appropriate subspace of the Hilbert space $\mathcal{H}$ of random variables is not related to the present discussion, because the underlying manifold is different. There is instead a clear connection with information geometry: to the best of the author’s knowledge, however, the identification of the underlying manifold as determined by the mathematical model of the experiment, as well as the identification of the map from the space $\mathcal{H}$ (where the modeling is abstractly done) to the space $\mathbb{R}^{n}$ (where the computations are carried out) have not yet been formalized.
5 Conclusions
@@
A mathematical model can properly be defined as an operator yielding a random variable $\boldsymbol{X}$ approximating the objective random variable $\boldsymbol{X}^{o}$ in $\mathcal{H}$. The outcome $\omega$ maps $\mathcal{H}$ onto case space $\mathbb{R}^{n}$. Estimation for linear models in case space consists in projecting the observed value $\boldsymbol{x}^{o}$ onto the prediction surface $\mathcal{M}_{\boldsymbol{x}}$}, and corresponds via $\omega$ to projecting $\boldsymbol{X}^{o}$ onto the model manifold $\mathcal{M}_{\boldsymbol{X}}$}. For nonlinear models a local optimum $\hat{\boldsymbol{x}}\in{\mathcal{M}_{\boldsymbol{x}}}$ can be found numerically, and is the image under $\omega$ of a locally optimal random variable $\hat{\boldsymbol{X}}\in{\mathcal{M}_{\boldsymbol{X}}}$. Asymptotic confidence regions can be computed on $\mathcal{M}_{\boldsymbol{x}}$} for linear models. For nonlinear models approximate asymptotic confidence regions can be computed on the affine tangent space $\mathcal{T}_{\hat{\boldsymbol{x}}}$ (constructed via computation of the Jacobian at the local optimum), which corresponds to the theoretical tangent space $\mathcal{T}_{\hat{\boldsymbol{X}}}\subset\mathcal{H}$.